Solve the differential equation
(x-y^2)dx+2xy=0
1
Expert's answer
2020-11-25T15:26:13-0500
"\begin{aligned}
(x-y^2)dx + 2xy dy &= 0\\
2y\frac{dy}{dx} - \frac{y^2}{x} &= -1\\
\\
Let\ v = y^2 \\
dv = 2ydy\\
\\
Then, \frac{dv}{dx} - \frac{v}{x} &= -1\ ---(i)\\
\\
P &= \frac{-1}{x}\\
\int{Pdx} &= -\int{\frac{1}{x}dx} = -\ln x \\ &= \ln{\frac{1}{x}}\\
\\
\therefore The\ I.F &= e^{ln\frac{1}{x}} = \frac{1}{x}


\end{aligned}"

Multiplying equation (i) by the I.F, we get,

"\begin{aligned}
\frac{1}{x}\frac{dv}{dx} - \frac{1}{x}\frac{v}{x} &= -\frac{1}{x}\\

\frac{1}{x}\frac{dv}{dx} - \frac{v}{x^2} &= -\frac{1}{x}
\end{aligned}"


Integrating both sides with respect to x,

"\begin{aligned}
\frac{v}{x} &= -\int{\frac{1}{x}dx} + b = -\ln{x} + b\\
\\
\therefore y^2 &=- x\ln{x}+bx
\end{aligned}"